Hahn Company uses a job-order costing system. Its plantwide predetermined overhead rate uses direct labor-hours as the allocation base. The company pays its direct laborers $13.00 per hour. During the year, the company started and completed only two jobs—Job Alpha, which used 64,200 direct labor-hours, and Job Omega. The job cost sheets for the these two jobs are shown below: Job Alpha Direct materials ? Direct labor ? Manufacturing overhead applied ? Total job cost $ 1,620,000 Job Omega Direct materials $ 225,000 Direct labor 357,500 Manufacturing overhead applied 165,000 Total job cost $ 747,500 Required: 1. Calculate the plantwide predetermined overhead rate. 2. Complete the job cost sheet for Job Alpha.
Answer #1
Answer
Job hours used in job omega is 27500 ( 357500/13)
Predetermined overhead =6 (165000/27500)
Labour cost = 64200*13 = 834,600
Manufacturing overhead = 64200*6 = 385,200
Direct materials = total cost - labour cost - manufacturing overhead
Direct material = 1620000-834,600-385200 = $ 400,600
Job alpha
Direct material $        400,600
Direct labour $        834,600
Manufacturing oh $        385,200
Tptal job cost $     1,620,400

Please like